2. Suppose a bond pays $1000 two years from now. The current market rate is 4% on the saving account. How much will pay for this if you purchase this bond today?
Answer #1

Given

Discount rate=i=4%

Time period=n=2 years

Future value of deposit=FV=$1000

Price of bond=Present value of future cash flow=FV/(1+i)n

Price of bond=1000/(1+4%)2=$924.56